Suppose you toss a coin 100 times and get 66 heads and 34 tails. Based on these results, what is the probability that the next flip resuming a tail



1
Expert's answer
2022-02-08T16:59:51-0500

Let the number of tails in the previous trials be "x=34"

Let the total number of trials be "n=100"

The number of heads in the previous trials is "n-x=100-34=66"

The probability that the next flip will be a tail is given as,

"p(tail)={x\over n}={34\over100}=0.34"